OpenStudy (anonymous):

Using the formula r= d over t where d is the distance in miles, r is the rate and t is the time in hours, at which rate must you travel to cover 227.5 miles in 3.5 hours?

OpenStudy (anonymous):

still want the answer? o; lol

OpenStudy (anonymous):

uh... yeah...

OpenStudy (anonymous):

r=d/t where d= 227.5 and t=3.5 so, r=227.5/3.5